Let R be the relation on the set A = {a, b, c, d, e, f} and
R = {(a,c), (b,d), (c,a), (c,e), (d,b), (d,f), (e,c), (f,d)}
Find the transitive closure pf R using Warshall’s algorithm.
1
Expert's answer
2020-10-29T20:32:00-0400

"R = \{(a,c), (b,d), (c,a), (c,e), (d,b), (d,f), (e,c), (f,d)\}"


Here are the steps of the Warshall’s algorithm:


Step 1. Assign initial values "W=M_R, k=0".


Step 2. Execute "k:=k+1."


Step 3. For all "i\ne k" such that "w_{ik}=1", and for "j" execute the operation "w_{ij}=w_{ij}\lor w_{kj}."


Step 4. If "k=n", then stop: we have the solution "W=M_{R^*}", else go to the step 2.



"n=|A|=6."


"W^{(0)}=M_R=\left(\begin{array} {cccccc} 0 & 0 & 1 & 0 & 0 & 0\\ 0 & 0 & 0 & 1 & 0 & 0\\ 1 & 0 & 0 & 0 & 1 & 0 \\ 0 & 1 & 0 & 0 & 0 & 1 \\ 0 & 0 & 1 & 0 & 0 & 0 \\ 0 & 0 & 0 & 1 & 0 & 0\end{array}\right)"



"W^{(1)}=\left(\begin{array} {cccccc} 0 & 0 & 1 & 0 & 0 & 0\\ 0 & 0 & 0 & 1 & 0 & 0\\ 1 & 0 & 1 & 0 & 1 & 0 \\ 0 & 1 & 0 & 0 & 0 & 1 \\ 0 & 0 & 1 & 0 & 0 & 0 \\ 0 & 0 & 0 & 1 & 0 & 0\end{array}\right)"



"W^{(2)}=\left(\begin{array} {cccccc} 0 & 0 & 1 & 0 & 0 & 0\\ 0 & 0 & 0 & 1 & 0 & 0\\ 1 & 0 & 1 & 0 & 1 & 0 \\ 0 & 1 & 0 & 1 & 0 & 1 \\ 0 & 0 & 1 & 0 & 0 & 0 \\ 0 & 0 & 0 & 1 & 0 & 0\end{array}\right)"



"W^{(3)}=\left(\begin{array} {cccccc} 1 & 0 & 1 & 0 & 1 & 0\\ 0 & 0 & 0 & 1 & 0 & 0\\ 1 & 0 & 1 & 0 & 1 & 0 \\ 0 & 1 & 0 & 1 & 0 & 1 \\ 1 & 0 & 1 & 0 & 1 & 0 \\ 0 & 0 & 0 & 1 & 0 & 0\end{array}\right)"


"W^{(4)}=\left(\begin{array} {cccccc} 1 & 0 & 1 & 0 & 1 & 0\\ 0 & 1 & 0 & 1 & 0 & 1\\ 1 & 0 & 1 & 0 & 1 & 0 \\ 0 & 1 & 0 & 1 & 0 & 1 \\ 1 & 0 & 1 & 0 & 1 & 0 \\ 0 & 1 & 0 & 1 & 0 & 1\end{array}\right)"



"W^{(5)}=\left(\begin{array} {cccccc} 1 & 0 & 1 & 0 & 1 & 0\\ 0 & 1 & 0 & 1 & 0 & 1\\ 1 & 0 & 1 & 0 & 1 & 0 \\ 0 & 1 & 0 & 1 & 0 & 1 \\ 1 & 0 & 1 & 0 & 1 & 0 \\ 0 & 1 & 0 & 1 & 0 & 1\end{array}\right)"



"M_{R^*}=W^{(6)}=\left(\begin{array} {cccccc} 1 & 0 & 1 & 0 & 1 & 0\\ 0 & 1 & 0 & 1 & 0 & 1\\ 1 & 0 & 1 & 0 & 1 & 0 \\ 0 & 1 & 0 & 1 & 0 & 1 \\ 1 & 0 & 1 & 0 & 1 & 0 \\ 0 & 1 & 0 & 1 & 0 & 1\end{array}\right)"


Therefore, the transitive closure of "R" is the following:


"R^* = \{(a,a),(a,c),(a,e), (b,b),(b,d),(b,f), (c,a),(c,c), (c,e), (d,b),(d,d),"

"(d,f),(e,a), (e,c),(e,e),(f,b), (f,d),(f,f) \}"